# “Here’s Your Evidence”: False Consensus in Public Twitter Discussions of COVID-19 Science

### v1

@@ 7.3 Theme 2: Discrediting and Diminishing
Case 7. This paper [64] examines HCQ as a COVID-19 prophylactic and finds no significant reduction in COVID-19 incidence. Although categorized as conformist-consistent, it mostly attracts contrarians for two reasons. First, it reports a positive but statistically non-significant effect of HCQ. Second, it acknowledges funding from the Minnesota Chinese Chamber of Commerce (MCCC), which contrarians use to question the study despite a clarification that funders played no role in conducting the study.
@@
Case 8. This study [33], classified as contrarian-consistent, finds that HCQ is effective against COVID-19, especially if used alongside azithromycin. Many think that it has initiated the debate around HCQ for COVID-19, as Donald Trump referred to it while still serving as US president [81].
@@ 7.4 Theme 3: Focus on Contextual Factors
In one case, conformist users draw attention to a contrarian-consistent paper [49] while focusing on contextual factors surrounding it. The study analyzes HCQ’s safety and reports short-term safety when prescribed alone but elevated cardiovascular risks when prescribed with azithromycin. The paper does not “position” itself as contrarian per se, but we assign a contrarian-consistent label due to the HCQ safety finding.
@@ 7.5 Theme 4: Dissemination of Oppositional Findings
Case 10. This paper [18], released in April 2020, shows that HCQ reduces recovery time and promotes pneumonia absorption. It is mostly discussed by conformists with mixed reception. Some tweets highlight study limitations like sample size and lack of clarity behind methodological choices, with some even insinuating that its conclusions are unreliable. Others express hope that HCQ may be a promising drug in treating COVID-19, though such expressions are reserved, often with calls for further evidence.
@@
Case 11. This study [31] concludes that variants may escape mRNA vaccine neutralization. Conformist users mostly circulate or summarize the study and discuss the concerning findings around immunization escape. One conformist science journalist suggests that the paper raises deep concerns about vaccines but receives criticism from others for being unduly “alarmist.” One of the authors also responds to the journalist, urging caution in interpreting results as the study is an in-vitro examination of a pseudovirus, not a direct assessment of vaccine effectiveness. Generally, conformists do not criticize the study and instead share it for informational purposes.
@@
Case 12. This study [30] of six vaccinated people infected with COVID-19 concludes that the Delta variant may possess vaccine evasion properties. It is predominantly discussed by conformists on Twitter.
@@ 8 Discussion and Conclusion
Next, we studied the temporal evolution of this conversation, activity, and new accounts around the declaration of a global pandemic (RQ3). Contrarian users became more active during the pandemic, both in terms of tweet volume and new account creation. Moreover, we found a collaborative effort to amplify prominent contrarian voices, even though we could not verify if this activity was organic or coordinated. Finally, we examined whether/how COVID-19 papers are misrepresented (RQ4). While not the norm, misinformation occasionally surfaces and garners attention. Some users, particularly contrarians, selectively present findings to bolster their views or question opposing research. Nevertheless, conformist users often engage in nuanced methodological discussions and share findings beneficial for the COVID-19 response, regardless of alignment with their stance.
@@ 8.1 Implications
Public sentiment should ideally reflect scientific consensus as closely as possible for topics where this consensus is the best proxy for objective truth (e.g., COVID-19 and climate change). False consensus allows duplicitous doubting of the evidence [59] and, consequently, the entertainment or even adoption of empirically flawed policies. For example, the very misperceptions we demonstrate have been used as reasoning for anti-scientific and potentially harmful policy positions by Florida’s Governor and Surgeon General [11]. Our findings highlight the need to quantify true scientific consensus on topics for which public and policy opinion deviate from the extant evidence. This is an open challenge for HCI and CSCW communities, as several systems need to be implemented for this to happen. These range from dynamic estimations of scientific consensus as it evolves, to implementing methods of digital delivery on educating users and broader networks on this consensus.
Reliance on official institutions as messengers may be insufficient for two reasons. First, trust in these institutions cannot always be assumed [45]. Second, in their role as aggregators of consensus, institutions may inadvertently reverse perceptions of such consensus. Our study highlights how many secondary sources cite a small number of primary sources to create an illusion of consensus; this can, in turn, give credibility to ideas without sufficient evidence to support them [84]. With the institutional approach, the opposite occurs; a large number of primary sources are narrowed through the lens of only a handful of secondary sources that are not personified and, thus, perhaps not as accountable in the eyes of the public.
@@ 8.2 Limitations
Naturally, our work is not without limitations. Although our abstract-count approach is similar to prior work [22], we acknowledge that paradigm shifts or scientific opinion switches can occur owing to authoritatively robust and large studies. To that end, another potential method of estimating scientific consensus would be a direct survey of experts who have published work in a given field or do active research within it. Nonetheless, we have no reason to believe that our main conclusions would change had we followed this approach, especially since we were fairly conservative in developing our codebooks. On the contrary, given findings like early hope for HCQ as a treatment drug followed by subsequent convergence against it, we expect that direct surveying of experts would only amplify our results. Overall, our case studies reveal that the authors of the studies we annotate tend to agree with our labels; cases where this is not as obvious were labeled as contrarian-consistent, which further supports that the false consensus effect we find may be even more pronounced in reality.
@@
Furthermore, we do not appraise methodological robustness or confidence in each study’s findings. However, we stress that we only aim to study false consensus; that is, the deviation between the number of primary and secondary sources behind certain claims. This is also why we explicitly study tweets that directly link to these primary sources. Quality control would be best situated within the scope of studies that appraise epistemic certainty itself, for example, meta-analyses or systematic reviews. Regardless, our cursory observations and quality-proxy analyses (e.g., publication odds) suggest that quality appraisal would not alter our main conclusions.
@@
Finally, the false consensus effect we report is derived from Twitter posts. This does not necessarily mean that the percentages we report reflect the opinions or attitudes of the entirety of Twitter, let alone the general population. For example, misinformed users may simply be more likely to share posts than others [61]; without a fully random sample, it is difficult to estimate how well this false consensus effect generalizes. Still, our findings reflect what may be directly perceivable on social platforms like Twitter, which means that opinions and behaviors may still be unduly influenced owing to descriptive social norms [19].
8.3 Ethical Statement and Disclaimer
This study received ethical approval from UCL Ethics Review Board. We only collect and use the minimal amount of data required to perform our analyses as per our research questions and follow-up analyses, and avoid direct user quotations to protect their identities.
Disclaimer. Although we use the terms “contrarian” and “conformist” to generally refer to users that are against or for scientific consensus, respectively, we do not wish to imply the ‘correctness’ of their respective views. We only study how science is used holistically among public circles and whether reasonable inferences are drawn from such discussions. For example, it would not be wrong to say that COVID-19 vaccines may lose some efficacy against newer variants. However, it is inaccurate to extrapolate this assertion to claims that COVID-19 vaccines are generally ineffective. “Contrarians” and “conformists” are merely a distinction based on how accurately they portray the state of scientific discourse.
Similarly, we are cautious not to describe scientific papers as contrarian or conformist; rather, we use the terms “contrarian-consistent” and “conformist-consistent.” In this way, we wish to convey that we classify papers based on whether their conclusions align with specific viewpoints, not on whether papers themselves have a certain “agenda.” Indeed, such a practice would be counter-intuitive. Studying and documenting any risks associated with officially espoused measures, among other things, is crucial to forming robust, evidence-based policy. Challenging conventional wisdom is pivotal in scientific advancement, and labeling science that brings new evidence to bear as “contrarian” would be a misnomer. To this end, we also clarify that we are neither assuming nor implying any authors’ stances or attitudes in classifying preprints. Put simply, we are not raising any issues regarding how COVID-19 science is conducted; this is beyond the scope of our work. We are only studying how this science is disseminated and discussed to ensure that it has the intended impact.
Appendix
@@ Appendix A Twitter Engagement Analyses
[…]
@@
Figure 10: Median tweets per day by account creation date per group.
(a) Lorenz curves of out-degrees per group and period.
(b) Lorenz curves of in-degrees per group and period.
Figure 11: Lorenz curve plots of network degree inequality.
@@ Appendix B Paper Publication Odds
[…]

### v2

@@ 7.3 Theme 2: Discrediting and Diminishing
Insinuations of financial conflicts. Case 7 [71] (later published in Clinical Infectious Diseases) examines HCQ as a COVID-19 prophylactic and finds no significant reduction in COVID-19 incidence. Although categorized as conformist-consistent, it mostly attracts contrarians for two reasons. First, it reports a positive but statistically non-significant effect of HCQ. Second, it acknowledges funding from the Minnesota Chinese Chamber of Commerce (MCCC), which contrarians use to question the study despite a clarification that funders played no role in conducting the study.
@@
Discussions of methodological limitations. Case 8 [35] (later published in the International Journal of Antimicrobial Agents), classified as contrarian-consistent, finds that HCQ is effective against COVID-19, especially if used alongside azithromycin. Many think that it has initiated the debate around HCQ for COVID-19, as Donald Trump referred to it while still serving as US president [91].
@@ 7.4 Theme 3: Focus on Contextual Factors
In Case 9 [53] (later published in The Lancet Rheumatology), conformist users draw attention to a contrarian-consistent paper while focusing on contextual factors surrounding it. The study analyzes HCQ’s safety and reports short-term safety when prescribed alone but elevated cardiovascular risks when prescribed with azithromycin. The paper does not “position” itself as contrarian per se, but we assign a contrarian-consistent label due to the HCQ safety finding.
@@ 7.5 Theme 4: Dissemination of Oppositional Findings
Hope in pre-consensus evidence. Case 10 [19], released in April 2020 before a broader consensus emerged, shows that HCQ reduces recovery time and promotes pneumonia absorption. It is mostly discussed by conformists with mixed reception. Some tweets highlight study limitations like sample size and lack of clarity behind methodological choices, with some even insinuating that its conclusions are unreliable. Others express hope that HCQ may be a promising drug in treating COVID-19, though such expressions are reserved, often with calls for further evidence.
@@
Clarifying scientific implications. Case 11 [33] (later published in Cell) concludes that variants may escape mRNA vaccine neutralization. Conformist users mostly circulate or summarize the study and discuss the concerning findings around immunization escape. One conformist science journalist suggests that the paper raises deep concerns about vaccines but receives criticism from others for being unduly “alarmist.” One of the authors also responds to the journalist, urging caution in interpreting results as the study is an in-vitro examination of a pseudovirus, not a direct assessment of vaccine effectiveness. Generally, conformists do not criticize the study and instead share it for informational purposes.
@@
Information-sharing and urging vigilance. Case 12 [32] (later published in BMC Medicine) studies six vaccinated people infected with COVID-19 and concludes that the Delta variant may possess vaccine evasion properties. It is predominantly discussed by conformists on Twitter.
@@ 8 Discussion and Conclusion
Next, we studied the temporal evolution of this conversation, activity, and new accounts around the declaration of a global pandemic (RQ3). Contrarian users became more active during the pandemic, both in terms of tweet volume and new account creation. Moreover, we found a collaborative effort to amplify prominent contrarian voices, even though we could not verify if this activity was organic or coordinated. Finally, we examined whether/how COVID-19 papers are misrepresented (RQ4). While not the norm, misinformation about the contents of studies occasionally surfaces and garners attention. Some users, particularly contrarians, selectively present findings to bolster their views or question opposing research. Nevertheless, conformist users often engage in nuanced methodological discussions and share findings beneficial for the COVID-19 response, regardless of alignment with their stance.
@@ 8.1 Implications
Public sentiment should ideally reflect scientific consensus as closely as possible for topics where this consensus is the best proxy for objective truth (e.g., COVID-19 and climate change). False consensus allows duplicitous doubting of the evidence [66] and, consequently, the entertainment or even adoption of empirically flawed policies. For example, the very misperceptions we demonstrate have been used as reasoning for anti-scientific and potentially harmful policy positions by Florida’s Governor and Surgeon General [11]. Our findings highlight the need to quantify true scientific consensus on topics for which public and policy opinion deviate from the extant evidence.
This is an open challenge for HCI and CSCW communities, as existing CSCW frameworks can lend themselves to addressing this problem in practice. Misinformation, both about COVID-19 specifically (e.g., [12, 84]) and on general issues more broadly (e.g., [26, 29, 44, 59, 61]), has received a lot of attention within the CSCW community especially with respect to how it is situated between social and technological systems. Due to challenges like lower trust in institutions that attempt to curtail problematic content [12] and the rapid rate at which misinformation proliferates, there have been recent calls to adopt approaches that target the social norms that enable the spread of misinformation, rather than individual pieces of information themselves [3]. Indeed, convergence towards such perspectives arises from long-standing work in the field that treats information as latent narratives [64, 87, 88]. Adherence to misinformation may then be understood as deep-rooted motivations, such as the need to protect one’s social identity or the need to resolve ambiguity around complex events [29], which are served by these narratives.
The idea that science around COVID-19 does not align with official guidelines is, itself, such a narrative. Crucially, belief in this narrative could then give rise to more layered beliefs, such as the idea that powerful institutions like social media platforms are censoring the truth [87]; something which our findings refute, since contrarian viewpoints do not reflect the extant evidence, nor are they censored based on the prominence of these views on Twitter. Successfully challenging these narratives to minimize reactance is something that can be informed by design perspectives within CSCW. For example, some works have proposed how subtle nudges may be used to address cognitive biases during information parsing like targeting the balance of availability between misinformation and corrections [50]. Other researchers have put forward recommendations for designing tools in science communication specifically, like providing analytics that enable science communicators to assess the audiences they reach and the impact of their efforts [86]. These design insights may be adapted to make scientific consensus more easily perceptible by the public, ranging from dynamic estimations of scientific consensus as it evolves, to implementing methods of digital delivery on educating users and broader networks on this consensus. The combination of misinformation and design frameworks can enable CSCW scholars to more effectively challenge complex and deep-seated latent beliefs, potentially introducing dissonance in people who adhere to these narratives and assisting them in forming more accurate perceptions [29].
Reliance on official institutions alone as messengers may be insufficient for two reasons. First, trust in these institutions cannot always be assumed [48]. Second, in their role as aggregators of consensus, institutions may inadvertently reverse perceptions of such consensus. Our study highlights how many secondary sources cite a small number of primary sources to create an illusion of consensus; this can, in turn, give credibility to ideas without sufficient evidence to support them [94]. With the institutional approach, the opposite occurs; a large number of primary sources are narrowed through the lens of only a handful of secondary sources that are not personified and, thus, perhaps not as accountable in the eyes of the public.
@@ 8.2 Limitations
Naturally, our work is not without limitations. Although our abstract-count approach is similar to prior work [23], we acknowledge that paradigm shifts or scientific opinion switches can occur owing to authoritatively robust and large studies. Indeed, particularly with respect to recommended pharmaceuticals for the treatment of COVID-19, we find that some official guidelines changed throughout our observation period. We do not consider these shifts in annotating papers (i.e., we only consider official guidelines at the time of annotation), since papers that could be considered consistent with official guidelines at the time may continue to be mentioned on Twitter after the guidelines change. However, we recognize limitations with this approach, as this may inflate the percentage of papers that we annotate as contrarian-consistent. For example, some papers put forward preliminary findings supporting the use of HCQ which was under emergency use authorization for the treatment of COVID-19 by the US Food and Drug Administration (FDA) until June 15th, 2020, when this authorization was revoked in the face of new evidence. [^4] https://bit.ly/3QH6R1C To that end, another potential method of estimating scientific consensus would be a direct survey of experts who have published work in a given field or do active research within it. Nonetheless, we have no reason to believe that our main conclusions would change had we followed this approach, especially since we were fairly conservative in developing our codebooks. On the contrary, given findings like early hope for HCQ as a treatment drug followed by subsequent convergence against it, we expect that direct surveying of experts would only amplify our results. Overall, our case studies reveal that the authors of the studies we annotate tend to agree with our labels; cases where this is not as obvious were labeled as contrarian-consistent, which further supports that the false consensus effect we find may be even more pronounced in reality.
@@
Furthermore, we do not appraise methodological robustness or confidence in each study’s findings. However, we stress that we only aim to study false consensus and illusory consensus; that is, the deviation between the number of primary and secondary sources behind certain claims, and how that may affect perceived norms on Twitter. This is also why we explicitly study tweets that directly link to these primary sources. Quality control would be best situated within the scope of studies that appraise epistemic certainty itself, for example, meta-analyses or systematic reviews. Regardless, our cursory observations and quality-proxy analyses (e.g., publication odds) suggest that quality appraisal would not alter our main conclusions.
@@
Finally, the false consensus effect we report is derived from Twitter posts. This does not necessarily mean that the percentages we report reflect the opinions or attitudes of the entirety of Twitter, let alone the general population. For example, misinformed users may simply be more likely to share posts than others [68]; without a fully random sample, it is difficult to estimate how well this false consensus effect generalizes. We also note that, while the behavior of some accounts suggests automated or at least inorganic activity, we do not eliminate these accounts from our analyses for two major reasons. First, due to changes to the platform’s API, we were unable to verify whether these accounts are indeed coordinated. Second, these accounts can act as information brokers for other organic users, such that removing them would obfuscate the nature of COVID-19 science discussions and their spread. Still, our findings reflect what may be directly perceivable on social platforms like Twitter, which means that opinions and behaviors may still be unduly influenced owing to descriptive social norms [20].
Acknowledgements. This material is based upon work supported by the National Science Foundation under Grant No. IIS-2046590, the UK’s EPSRC grant EP/S022503/1, which supports the UCL Centre for Doctoral Training in Cybersecurity, and the UK’s National Research Centre on Privacy, Harm Reduction, and Adversarial Influence Online (REPHRAIN, UKRI grant: EP/V011189/1). Any opinions, findings, and conclusions or recommendations expressed in this work are those of the authors and do not necessarily reflect the views of the NSF, UKRI, or EPSRC.
@@ Appendix A Twitter Engagement Analyses
Beyond Twitter engagement at the paper level, we also assess whether a user’s or tweet’s stance is associated with Twitter engagement. We compute the total retweets that each user accumulates on their COVID-19 paper tweets and regress this on user stance while controlling for number of followers and accounts they are following, total tweets posted by the user, number of tweets that appear in our dataset, user verification status, and account age (in days). The model is significant and explains a moderate degree of variance ($\textit{F}_{(7,81529)}=18.25,\textit{p}<0.001,R^{2}=0.379$), but only the number of followers significantly predicts total retweets accumulated ($\beta=0.62,\textit{p}=0.02$). User stance does not show significant effects ($\beta=-0.003,\textit{p}=0.55$). All continuous variables are z-score standardized, and we employ robust standard errors.
At the (English-language) tweet level, however, we observe a slightly different pattern. We regress the retweets obtained by individual tweets on whether a contrarian or conformist user posts the tweet. We control for all of the account features used in the user-level analysis while additionally controlling for the age of the tweet itself, and whether the tweet is a reply, a quote tweet, or original; again, we employ z-score standardization and robust standard errors. This results in a significant overall model ($\textit{F}_{(9,360621)}=149.5,\textit{p}<0.001,R^{2}=0.118$), though only 11.8% of the variance in the outcome variable is explained. We show the full regression in Table 4.
@@ Appendix B Paper Publication Odds
The model is statistically significant (p $<0.001$), but the overall fit is low (pseudo-$R^{2}=0.03$). As expected, the number of authors ($\beta=0.35,\textit{p}<0.01$) and paper recency ($\beta=-0.39,\textit{p}<0.001$) show significant and moderate effects. Paper stance does not statistically influence publication odds ($\beta=-0.20,\textit{p}=0.40$), though we note the non-negligible effect size. Thus, owing to the relatively low number of contrarian-consistent papers, this finding may be inconclusive due to potentially limited statistical power in detecting a true effect.
@@ C.1 Suspicious Account Activity
Figure 10: Median tweets per day by account creation date per group.
@@ C.2 Suspicious Account Amplification
(a) Lorenz curves of out-degrees per group and period.
(b) Lorenz curves of in-degrees per group and period.
Figure 11: Lorenz curve plots of network degree inequality.